OpenStudy (anonymous):

Evaluate to three decimal places. e^1.35

OpenStudy (anonymous):

\[e ^{x}=1+\frac{ x }{ 1! }+\frac{ x ^{2} }{2! }+\frac{ x ^{3} }{3! }+...\]